550V is applied to a 2600-pF capacitor.


How much energy is stored?
Express your answer using three significant figures and include the appropriate units.
Answer #1

energy = 1/2*C*v^2


= 1/2* 2600*10^-12 * 550^2


= 3.93 * 10^-4 J
Answer #2

Capacitance=2600 pF=2600*10-12 F

Voltage applied=550V

Energy strored=0.5CV2

=0.5*2600*10-12*5502

=3.93*10-4 J
Answer #3

E= (1/ 2) * C * V ^2 = 0.5 * 2600 * 10 ^ -12 * 550 * 550 = 39325 * 10 -8 J